OpenStudy (anonymous):

HELP PLEASEEEE! F = 27.2 x m / t is the formula F is the force M is the mass T is the time in seconds Solve for M Please show your work f is 11,000 & t is 8. what i got for m was 1,375... am i correct ?

OpenStudy (anonymous):

Nope :-( \[\Large \begin{array}{l} F = \frac{{27.2m}}{t}\\ 27.2m = Ft\\ m = \frac{{Ft}}{{27.2}} \end{array}\]

OpenStudy (anonymous):

aww, well i tried! lol thank you :D

OpenStudy (anonymous):

Haha and that's the main thing! No worries =)

OpenStudy (anonymous):

so the answer would be m = 3,235.29 ?

OpenStudy (anonymous):

Yeah =)

OpenStudy (anonymous):

:D great, are you busy right now?